OpenStudy (anonymous):

can anyone please help me with trigonometry?? I've been asking and asking through messages and no one replies

OpenStudy (jdoe0001):

tis easier with an specific posting

OpenStudy (anonymous):

Describe how the graphs of functions f and g are related : f(x) = 2 cos(pi x) and g(x) = 2cos (2pi x)

OpenStudy (jdoe0001):

have you covered function transformations? like for say a quadratic function or a linear one?

OpenStudy (anonymous):

and second question : Determine the period,domain,range,zeros, and asymptotes (if any). Then sketch graph of function y = -3tan (1/2 x)

OpenStudy (anonymous):

yes well the basics is what I know for example y = Asin(B(x+/- c)) +/- D

OpenStudy (anonymous):

so can you help me with either question please?

OpenStudy (anonymous):

I've done about 15 other questions just need to finish those two

OpenStudy (anonymous):

oh and amplitude and period of y = -sin (x - pi/4) + 2

OpenStudy (anonymous):

that's it

OpenStudy (jdoe0001):

well, for one f(x) and g(x) share the same amplitude

OpenStudy (jdoe0001):

and for another, g(x) is the same as f(x), just SHRUNKEN a bit due to its phase/horizontal shift \(\bf f(x) = 2 cos({\color{blue}{ \pi }}x) \qquad g(x) = 2cos ({\color{blue}{ 2\pi}} x)\\ \quad \\ \textit{phase shift}\quad f(x)\implies \cfrac{2\pi}{{\color{blue}{ \pi }}}\qquad g(x)\implies \cfrac{2\pi}{{\color{blue}{ 2\pi}}}\)

OpenStudy (anonymous):

wait so there is a phase shift?

OpenStudy (anonymous):

I thought that was period when you do 2pi/ B

OpenStudy (jdoe0001):

ohh right, it's the period... .I meant... .thus is SHRUNKEN

OpenStudy (jdoe0001):

other than that... they're the same function

OpenStudy (anonymous):

so g(x) is shrunken version of f(x)?

OpenStudy (anonymous):

hi @ranga !

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

for how the graphs are related I have " Neither experience phase shift or vertical translation". "Both have same amplitude of 2." and "Period of f(x) = 2 and period of g(x) = 1. Anything else I should add?

OpenStudy (anonymous):

oh and g(x) is shrunken version of f(x). Wait horizontally or vertically?

OpenStudy (jdoe0001):

the period is being changed from g(x) so it'll shrink from